OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. x-intercept 8/13 and y-intercept 20/27

OpenStudy (anonymous):

\(\large\text{y = mx + b}\) where b = y-intercept (what y is when x=0) = 20/27 m=slope Average slope: \[m=\frac{\triangle y}{\triangle x}=\frac{y_2-y_1}{x_2-x_1}\] x-intercept (what x is when y=0), 8/13 Looks like this: